OpenStudy (ksaimouli):

find the slope at 9-x^2 at x=2

OpenStudy (anonymous):

do you know how to find the derivative?

OpenStudy (ksaimouli):

no but i should use the f(a+h)-f(a)/h

OpenStudy (ksaimouli):

formula

OpenStudy (ksaimouli):

can u just give the function i will do rest of it

OpenStudy (anonymous):

well, f(x) = 9-x^2 does that help?

OpenStudy (ksaimouli):

nope

OpenStudy (ksaimouli):

i know that

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} f(a+h)-f(a)/h\] f(a+h)=9-(a+h)^2 f(a)=9-a^2 does that help?

OpenStudy (ksaimouli):

i figured it out myself thx for ur support